OpenStudy (anonymous):

Coulomb help please? Balloon is rubbed with wool giving it a charge of -8.9x10^-14 C. What is the force between balloon and a metal sphere that is charged 25 C and is 2 km away?

OpenStudy (anonymous):

I tried setting up the equation but I think its terribly wrong. F=8.9* -8.9x10^-14/2,000 m ?

OpenStudy (anonymous):

Mostly there :) You set up the electric potential of the balloon at 2km \[V_{balloon} = k\frac{q_{balloon}}{r} \] To get the force between two charged objects you have \[ F_{coulomb} = k\frac{q_1q_2}{r^2} \]

OpenStudy (anonymous):

Ahhh! I always get confused! >.< I thought that r2 was always supposed to be converted into meters?

OpenStudy (anonymous):

You're right :) You do need to convert the "r" into meters because of the constant, K you put in front. K is in units of Newton*meters^2/Coulomb^2 \[ K \propto \frac{N \ m^2}{C^2}\] So to get units of Force - Newtons - you have to multiply it by Coulomb^2/m^2, so that's what you do! \[ \frac{q_1q_2}{r^2} \propto \frac{C*C}{m^2}\] If you leave r in km, your answer here will be off by a magnitude of 1x10^-6 :P

OpenStudy (anonymous):

Will my answer be 5.0x10^-9 N ? @AllTehMaffs

OpenStudy (anonymous):

Indeed!

OpenStudy (anonymous):

Thanks for the help! Much appreciated (: